# Fine-tuning an ECG Foundation Model to Predict Coronary CT Angiography Outcomes

### v1

@@ DISCUSSION
This study constructs an interpretable AI-ECG model for predicting severe stenosis and total occlusion outcomes in CCTA using ECG. The model is based on ECGFounder, which has been trained for predicting new-onset MI, and fine-tuned using transfer learning to address shortcomings in existing studies regarding the study population and predicted outcomes. The model’s predictive performance was systematically validated using ROC curves, subgroup analysis, and box plots. Furthermore, based on the incidence distribution of major coronary artery lesions, a prediction probability threshold was set to define high-risk groups, and the model’s risk stratification capability was further evaluated using calibration curves and cumulative event occurrence curves.
@@
The main innovations of this study are reflected in the following aspects: First, the proposed model can directly extract clinically relevant information on severe and complete CCTA blockage from the raw ECG signal, and still shows good consistency in ECGs that are clinically diagnosed as normal, suggesting the potential application value of the model in screening for occult CAD; Second, the study systematically evaluated the model’s performance in risk stratification, and the results show that it has the ability to assist in clinical decision stratification; Finally, through comparative analysis of ECG waveform characteristics between high-risk and low-risk populations, this study attempts to explore potential pathological differences at the signal level, providing a new perspective for the study of CAD mechanisms.
While this study demonstrates some innovation in model construction and clinical application exploration, several limitations remain. First, the data primarily came from two centers, and the sample size and population diversity need further expansion, potentially limiting the model’s generalization ability. Second, due to the low incidence of severe and complete LM obstruction, the AUC of LM was skewed in the external validation set and among individuals with normal ECG during model training and validation, suggesting that the results may be affected by sample imbalance and require careful interpretation.
In future research, we will incorporate large-scale data from more centers to further improve the model’s stability and generalization ability. Simultaneously, we will conduct more refined analyses of LM lesions, such as combining spatial localization information from anterior and posterior leads to explore ECG characteristics corresponding to different anatomical regions, thereby improving the model’s ability to identify specific coronary artery lesions.
@@ METHODS
This study aimed to construct an interpretable AI-ECG model for predicting CCTA results from ECG, thus providing a potential alternative for non-invasive risk assessment of CAD. The model was developed based on 2,595 CCTA examinations and 4,620 paired ECG data from 2,323 patients at Peking University People’s Hospital, China. External validation was performed on 135 CCTA examinations and 228 paired ECG data from 135 patients at the Second Hospital of Tianjin Medical University, China, to systematically evaluate the model’s generalization performance across different centers and populations.
@@
To assess the clinical applicability of the model, this study constructed a systematic performance evaluation framework, comprehensively validating the model from multiple dimensions, including discriminative ability, probabilistic reliability, and clinical application potential. First, ROC curves were plotted to evaluate the model’s discriminative performance in identifying severe and complete occlusion in CCTA. Box plots of the model’s predicted probabilities for different degrees of stenosis were then used to quantitatively analyze the model’s ability to distinguish the severity of coronary artery stenosis. Furthermore, based on the model’s predicted probabilities and the actual incidence of severe coronary artery occlusion, the study population was stratified into high-risk and low-risk groups. Simultaneously, calibration curves were used to systematically test the reliability of the model’s predicted probabilities, and cumulative event occurrence curves for different risk groups were plotted to evaluate the model’s stratified predictive ability for future coronary events.
@@
Furthermore, this study systematically evaluated the model’s classification performance across different age groups, sexes, and subgroups with varying ECG and CCTA examination time differences. We further explored its ability to identify severe coronary artery stenosis on normal ECGs, thereby validating the model’s potential application value in screening for occult lesions in coronary artery disease (CAD). This evaluation process extends from classification performance to probabilistic reliability, and further to clinical decision support potential, constructing a complete and coherent clinical performance evaluation pathway for the model.
@@
To enhance the interpretability of the model, this study further compared the differences in single heartbeat waveforms in the twelve leads among different risk groups classified by the model, analyzed the key waveform regions that the model focuses on and their association with coronary artery lesions, thereby exploring the model’s discrimination criteria from an electrophysiological perspective, providing intuitive evidence for understanding the mechanism of AI-ECG in CAD identification, and providing a new research perspective for related clinical studies.
@@
This study included 2,595 CCTA examinations and 4,620 paired electrocardiogram (ECG) data from 2,323 patients at Peking University People’s Hospital in China for model development, and external validation was performed on 135 CCTA examinations and 228 paired ECG data from 135 patients at the Second Affiliated Hospital of Tianjin Medical University in China. To avoid bias caused by potential ECG bias due to clinical interventions such as catheterization lab procedures following CCTA examinations, the ECG examination time was strictly limited to be earlier than the corresponding CCTA examination time during data cleaning. This study has been approved by the Ethics Committee of Peking University (Approval No.: IRB00001052-23071), and the Ethics Committee has approved the waiver of the informed consent form.
Based on CCTA examination reports, under the guidance of professional clinicians, the original diagnostic text was structured, cleaned, and standardized using a large language model to extract stenosis information for the RCA, LAD, LCX, and LM. Based on clinical consensus, "normal, mild stenosis, and moderate stenosis" were categorized into the non-severe stenosis group, while "severe stenosis and complete occlusion" were categorized into the severe lesion group. This stratification method aligns with clinical criteria for determining lesions with significant prognostic value.
Meanwhile, the ECG diagnostic texts were uniformly cleaned and recoded, and the study population was further divided into two categories: "normal ECG" and "abnormal ECG" to support subsequent stratified analysis of the ability to identify occult coronary artery lesions.
@@
Furthermore, we conducted subgroup analyses on the internal validation set, stratifying the population based on age (<65<65 years vs. $\geq 65$ years), sex, and the time interval between ECG and CCTA examinations ($\leq 3$ hours vs. >3>3 hours). We systematically compared the differences in model performance among different subgroups to assess the model’s robustness under various population characteristics and time conditions. Finally, for patients with normal ECG diagnoses, we further analyzed the model’s prediction results to validate the potential application value in detecting occult CAD.
@@
Interpretability Analysis
The study stratifies patients according to predicted probabilities in the model and further compares the characteristic differences in the ECG waveforms between different risk groups to improve the interpretability of the AI-ECG model and reveal the electrophysiological basis of its risk assessment.
@@
ACKNOWLEDGMENTS
Shenda Hong is supported by the National Natural Science Foundation of China (62102008, 62172018), CCF-Tencent Rhino-Bird Open Research Fund (CCF-Tencent RAGR20250108), CCF-Zhipu Large Model Innovation Fund (CCF-Zhipu202414), PKU-OPPO Fund (BO202301, BO202503), Research Project of Peking University in the State Key Laboratory of Vascular Homeostasis and Remodeling (2025-SKLVHR-YCTS-02).
DECLARATION OF INTERESTS
The authors declare no competing interests.

### v2

@@ Discussion
This study constructed and validated an ECG-based artificial intelligence model for predicting severe stenosis and total occlusion lesions shown by CCTA. Based on the large-scale pre-trained ECG foundation model ECGFounder, the model was further adapted to coronary artery disease-related tasks through transfer learning strategies, thereby expanding the application scope of existing AI-ECG studies in predicting outcome types and study populations.
The model’s discriminative performance was systematically evaluated using multi-dimensional methods, including receiver operating characteristic (ROC) curve analysis, subgroup analysis, and prediction probability distribution analysis, demonstrating stable performance under different population characteristics and testing conditions.
Furthermore, this study incorporated the epidemiological characteristics of different coronary artery lesions and used vessel-specific prediction probability thresholds to stratify the population at risk. The reliability and longitudinal discriminative power of the model’s predictions were evaluated through calibration analysis and Kaplan-Meier curves. In addition, decision curve analysis showed that, within a range of clinically meaningful risk thresholds, the model-based risk stratification strategy had a higher net clinical benefit than the "treatment-all" and "treatment-none" strategies. Sensitivity analyses for different threshold selections also indicated that the model’s performance remained stable within reasonable limits.
@@
The main innovations of this study are reflected in the following aspects. First, the proposed model directly extracts clinically relevant information on severe stenosis and complete occlusion defined by CCTA from raw ECG signals, while maintaining stable performance in ECGs that are clinically interpreted as normal, highlighting its potential value for screening occult coronary artery disease. Second, the study systematically evaluated the model’s ability for risk stratification, demonstrating its capacity to support clinical decision-making; in particular, decision curve analysis further demonstrates that model-assisted risk stratification can deliver net clinical benefit across a range of clinically relevant risk thresholds. Finally, by comparing ECG waveform characteristics between model-defined high-risk and low-risk populations, this study explored potential signal-level differences associated with coronary artery disease, providing a novel perspective for investigating its underlying electrophysiological mechanisms.
While this study demonstrates methodological innovation and explores the clinical applicability of AI-enabled ECG analysis for coronary artery disease, several limitations should be acknowledged.
First, the data were derived from two medical centers, and although multicenter validation was performed, further expansion in sample size and population diversity is necessary to strengthen generalizability across broader clinical settings.
Second, owing to the low incidence of severe and complete left main (LM) coronary obstruction, model performance for LM lesions was less stable in external validation and among individuals with clinically normal ECGs, suggesting that class imbalance may influence performance estimates and that these results should be interpreted with caution.
Third, the current study focuses on predicting CCTA-defined coronary stenosis from ECG signals, and the relationship between model performance and downstream clinical decision-making requires further investigation. In particular, how model-guided risk stratification may interact with different clinical management strategies remains to be explored through prospective studies.
Fourth, because the training and validation datasets excluded patients who had undergone percutaneous coronary intervention (PCI) or coronary artery bypass grafting (CABG), the applicability of the proposed model for evaluating in-stent restenosis or graft vessel occlusion after revascularization may be limited. In addition, the model showed reduced performance among individuals with a prior history of myocardial infarction, which may be attributable to persistent ECG alterations resulting from previous infarction that obscure lesion-specific electrophysiological patterns. Dedicated studies incorporating post-revascularization populations and patients with prior myocardial infarction are therefore warranted.
Fifth, the study population primarily reflects a real-world CCTA-referred cohort, in which patients with suspected acute myocardial infarction caused by complete coronary occlusion are typically triaged directly to invasive coronary angiography rather than CCTA. Consequently, detailed annotations distinguishing acute from chronic ischemic conditions were not consistently available, limiting the ability to differentiate electrophysiological patterns associated with acute versus chronic coronary occlusion.
Finally, although the explainable analyses provide insights into ECG regions contributing to model predictions, the clinical significance of subtle signal-level variations not readily discernible by human interpretation remains unclear. Moreover, current explainability approaches have inherent methodological limitations, underscoring the need for further methodological refinement.
In future research, we plan to incorporate large-scale data from additional centers to further enhance the model’s stability and generalizability. In parallel, we will conduct more refined analyses of LM coronary lesions, including the integration of spatial information from anterior and posterior ECG leads to better characterize electrocardiographic patterns associated with distinct anatomical regions and to improve lesion-specific identification. Importantly, to clarify the relationship between model performance and clinical decision-making, we are currently designing a prospective randomized controlled trial (RCT) to evaluate the impact of AI-ECG–guided risk stratification on downstream diagnostic strategies and clinical outcomes in real-world settings.
@@ Methods
This study aimed to construct an AI-ECG model with explainable analysis capabilities to predict coronary artery lesions shown by CCTA based on ECG, thus providing a potential auxiliary tool for non-invasive risk assessment of CAD. The model was developed based on 2595 CCTA examinations and 4620 paired ECG data from 2323 patients at Peking University People’s Hospital. To systematically evaluate the model’s generalization ability across different centers and populations, we further validated it externally on 2477 CCTA examinations and 2477 paired ECG data from 2477 patients at the Second Hospital of Tianjin Medical University. In addition, this study included a longitudinal follow-up cohort of 400 patients from the Second Hospital of Tianjin Medical University to assess the model’s prospective risk stratification ability in real-world clinical scenarios.
@@
To comprehensively evaluate the model’s clinical application potential, this study constructed a systematic performance evaluation framework, validating the model from multiple dimensions, including discriminative ability, predictive probability reliability, and clinical decision support value. First, ROC curves were plotted to evaluate the model’s discriminative performance in identifying severe stenosis and total occlusion lesions shown by CCTA. Subsequently, by comparing the distribution of the model’s predicted probabilities for different degrees of stenosis, the ability to differentiate the severity of coronary artery lesions was quantitatively analyzed. Based on this, the study subjects were divided into high-risk and low-risk groups by combining the model’s predicted probabilities with the actual incidence of severe lesions. Simultaneously, calibration curves were used to systematically assess the reliability of model-predicted probabilities, while Kaplan–Meier analyses with log-rank tests were performed to evaluate the model’s ability to stratify the risk of future MI events.
@@
Furthermore, this study systematically analyzed the model’s classification performance in subgroups based on age, sex, medical history, and the time interval between ECG and CCTA examinations. It further evaluated the model’s ability to identify severe coronary artery stenosis in individuals with clinically normal ECGs, thus validating its potential value in screening for occult CAD. To avoid bias caused by different assessment granularities, we reported ROC analysis results at both the CCTA examination level and the patient level. Furthermore, to evaluate the model’s application prospects in assisting clinical decision-making, this study plotted DCA curves and supplemented them with sensitivity analyses based on different risk thresholds. The aforementioned evaluation process covers multiple levels, from classification performance and probabilistic reliability to clinical decision support potential, constructing a complete and coherent path for clinical performance validation of the model.
@@
To enhance the explainability of the model results, this study further compared the differences in single-beat waveforms in 12-lead electrocardiograms across different risk groups identified by the model, analyzed key waveform regions emphasized by the model and their potential association with coronary artery lesions, and explored the model’s discriminative criteria from an electrophysiological perspective. This analysis provides intuitive evidence for understanding the working mechanism of artificial intelligence electrocardiography in coronary artery disease identification and offers a new research perspective for related clinical and methodological studies.
@@
This study included data from 2323 patients at Peking University People’s Hospital for model development, encompassing 2595 coronary CT angiography (CCTA) examinations and 4620 paired electrocardiogram (ECG) recordings. These patients were identified after applying predefined exclusion criteria, whereby individuals with evidence of prior or subsequent percutaneous coronary intervention (PCI) or coronary artery bypass grafting (CABG) related to the indexed CCTA were excluded. External validation was performed using data from 2477 patients at the Second Hospital of Tianjin Medical University, including 2477 CCTA examinations and 2477 paired ECG records, after applying the same exclusion criteria. In addition, a longitudinal follow-up cohort comprising 400 patients from the Second Hospital of Tianjin Medical University was constructed to evaluate the model’s risk stratification capability during clinical follow-up.
To further minimize the potential influence of coronary revascularization procedures on ECG characteristics, all ECG recordings were required to precede the corresponding CCTA examination. This combined temporal and clinical exclusion strategy was adopted to reduce bias introduced by procedure-related ECG alterations.
The study was approved by the Peking University Medical Ethics Committee (Approval No.: IRB00001052-23071), Peking University People’s hospital Ethics Committee (Approval No.: 2024PHB428-001), and the ethical review committee of The Second Hospital of Tianjin Medical University(Approval No.: KY2025K386). And the requirement for informed consent was waived in accordance with relevant ethical guidelines.
Based on CCTA examination reports, under the supervision of professional clinicians, a large language model was used to perform structured parsing, cleaning, and standardization of the original imaging diagnostic text to systematically extract stenosis information for the right coronary artery (RCA), left anterior descending artery (LAD), left circumflex artery (LCX), and left main coronary artery (LM). According to clinically accepted standards, luminal stenosis was categorized as normal (<25%), mild stenosis (25–50%), moderate stenosis (50–70%), severe stenosis ($\geq 70\%$), or complete occlusion. For model development and evaluation, lesions classified as normal, mild stenosis, or moderate stenosis were grouped as non-severe stenosis, whereas severe stenosis and complete occlusion were grouped as severe lesions. All label generation and cleaning processes underwent manual auditing and confirmation by multiple clinicians to ensure consistency and reliability. This stratification scheme is consistent with clinical practice for identifying coronary artery lesions with established prognostic significance.
Simultaneously, the ECG diagnostic texts were uniformly cleaned and recoded, and the study population was further divided into two categories: "normal ECG" and "abnormal ECG," to support subsequent stratified analysis of the ability to identify occult coronary artery lesions. All ECG classifications were audited by clinicians.
@@
Furthermore, we conducted subgroup analyses on the internal validation set, stratifying the population based on age (<65<65 years vs. $\geq 65$ years), sex, medical history, and the time interval between ECG and CCTA examinations ($\leq 3$ hours vs. >3>3 hours). We systematically compared the differences in model performance among different subgroups to assess the model’s robustness under various population characteristics and time conditions. Finally, for patients with normal ECG diagnoses, we further analyzed the model’s prediction results to validate the potential application value in detecting occult CAD. Furthermore, we analyzed the data on exam-level and patient-level data to verify the robustness of the model.
@@
Explainability Analysis
The study stratifies patients according to predicted probabilities in the model and further compares the characteristic differences in the ECG waveforms between different risk groups to improve the explainability of the AI-ECG model and reveal the electrophysiological basis of its risk assessment.
@@
Declaration statements
Data Availability
The data that support the findings of this study are not publicly available due to restrictions imposed by institutional ethics committees and data governance policies of the participating hospitals. Access to the data may be considered upon reasonable request to the corresponding author, subject to approval by the relevant ethics committees.
Code Availability
The code used for model development and evaluation in this study is publicly available at https://github.com/shunxio/AnyECG-CCTA.
Acknowledgements
Shenda Hong is supported by the National Natural Science Foundation of China (62102008, 62172018), the CCF–Tencent Rhino-Bird Open Research Fund (CCF-Tencent RAGR20250108), the CCF–Zhipu Large Model Innovation Fund (CCF-Zhipu202414), the PKU–OPPO Fund (BO202301, BO202503), and the Research Project of Peking University in the State Key Laboratory of Vascular Homeostasis and Remodeling (2025-SKLVHR-YCTS-02). Kangyin Chen is supported by the National Natural Science Foundation of China (82470527). The authors thank all collaborators and participating institutions for their support and contributions to this research.
Author Contributions
Yujie Xiao led the conceptualization of the study, designed the methodology, implemented the models, performed validation and formal analyses, and drafted the initial version of the manuscript. Qinghao Zhao contributed to the manuscript revision process, assisted in refining the explanation of results, and provided critical feedback on the study. Jian Liu, Hao Zhang and Tong Liu contributed to data acquisition, preprocessing, and curation. Gongzheng Tang, Zhuoran Kan, Deyun Zhang, Jun Li, Guangkun Nie, Xiaocheng Fang, Haoyu Wang, and Shun Huang contributed to data preparation and investigation.Kangyin Chen provided clinical expertise and contributed to data interpretation.Shenda Hong supervised the entire project, provided resources, guided study design and interpretation, served as the corresponding author, and takes full responsibility for the integrity of the work. All authors reviewed, revised, and approved the final manuscript.
Competing Interests
Shenda Hong is an Associate Editor of npj Digital Medicine. Shenda Hong was not involved in the journal’s review of, or decisions related to, this manuscript. The other authors declare no competing financial or non-financial interests.